A ribbon of length 5¼ m is cut into small pieces each of length ¾m. Number of pieces will be:
(a) 5 (b) 6 (c) 7 (d) 8

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (c) is correct here
First we have to convert the mixed fraction into improper fraction 51⁄4 = 21/4
Then, 21/4 ÷ 3⁄4

= (21/4)/ (3⁄4)
= (21/4) × (4/3)
= (21 × 4)/ (4 × 3)
= (7 × 1)/ (1 × 1)
=7


Thanks